OpenStudy (anonymous):

5^71+5^72+5^73is completely divided by

OpenStudy (anonymous):

5^71+5^72+5^73 =5^71 * (1+5 +5^2) =5^71 * (1+5 +25) =31*(5^71)

OpenStudy (anonymous):

then

OpenStudy (anonymous):

a)150 b)155 c)160 d)30

OpenStudy (anonymous):

see 155=31*5 so B

OpenStudy (anonymous):

5^71=5

OpenStudy (anonymous):

how u say 5^71 is 5

OpenStudy (anonymous):

see 5^71 =5*5^70 means 5 is factor or divides 5^71